OpenStudy (anonymous):

Julie drives 100 mi to Grandmother's house. On the way to Grandmother's, Julie drives half the distance at 44.0mph and half the distance at 72.0mph . On her return trip, she drives half the time at 44.0mph and half the time at 72.0mph . What is her average speed on the return trip?

OpenStudy (kropot72):

Let the time taken for the total return trip be t hours. Distance = speed * time Distance travelled at 44 mph = 44 * (t/2) ..........(1) Distance travelled at 72 mph = 72 * (t/2) ..........(2) The total distance travelled is 100 miles. Therefore \[100=(44\times\frac{t}{2})+(72\times\frac{t}{2})=22t+36t\ .......(3)\] From equation (3) we find that 58t = 100 Therefore the time for the return trip is \[t=\frac{100}{58}\ hours\ .........(4)\] The average speed for the return trip is given by \[Average\ speed=\frac{distance}{time}=\frac{100}{t}=\frac{100}{\frac{100}{58}}=\frac{100\times58}{100}\ mph\]

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."